# Attention (as Discrete-Time Markov) Chains

### v1

@@ A.4 More Segmentation Results
A.5 Illustration of Bouncing as Consolidation Mechanism
Figure 7 visualizes the first bounces and the steady state of two example images, heads and layers.
Figure 8 also illustrates that iterating over the Markov chain results in an iterative refinement of the map, consolidating meta-stable states. The steady state eventually visualizes the global incoming attention. This effect cannot be observed for an input image with random pixels (Figure 8, bottom). Instead, the Markov chain converges very quickly to a dispersed visualization after one step for a noisy input, as the attention matrix in not structured. We compute the visualizations by passing the depicted images through DINOv1 and extracting the attention map in the 9th layer and 3rd head. Then, we compute the first bounce ($n=1$) with a uniform initial vector, the 2nd bounce, and the steady state. The same effect can be seen in Figure 13, where the second bounce (Ours) is significantly cleaner and better adheres to the depicted objects compared to the first bounce (Column Select).
@@ A.5 Illustration of Bouncing as Consolidation Mechanism
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
A.6 Additional Results for Masking Most Influential Tokens
A.6.1 Per-Model Results
In Figure 9 we show the results of performing the most-influential token masking experiment described in the main paper for individual models. The area-under-curve (AUC) metric in the main paper is computed over the curves.
A.6.2 Per-Layer Results
In Figure 10, we show the result of choosing progressively more number of layers to mask for the most-influential token masking experiment conducted in the main paper. “6 (Ours)” indicates what we used in the paper. Layers are selected starting from the furthest from the output. The results are shown for DINOv2 (ViT-S/14) w/ registers over the same dataset used in the main paper. However, similar trends were observed for all architectures. We opted for masking tokens from only the first 50% of the layers to avoid distorting the output space and a too steep performance drop (which happens for $n>6$), while still having an effect on the classification results for higher token mask percentage (which does not occur for $n<6$).
A.7 $\lambda_{2}$ Weighting for Masking Most Influential Tokens Experiment
We performed another token masking experiment similar to the one conducted in of the paper, only using TokenRank for determining the masking order. In this experiment, we compare using uniform weights per head (as in main paper), random head weights, and using $\lambda_{2}$ weighting. Results can be seen in Figure 11. Perhaps unsurprisingly, randomly weighting the heads performs better than uniform weights, as the transformer is more sensitive to individual heads being fully or almost fully masked out, rather than gradually masking out all heads together. Importantly, results indicate that $\lambda_{2}$ weighting can be used for increasing the degradation in accuracy, suggesting it can serve as a useful tool for determining importance of heads. This was also observed to improve zero-shot segmentation results (see Section A.2).
Linear Probing Experiment
Additionally, we also train a linear classifier on top of the TokenRank vectors (incoming attention) per layer, where we use different weighting schemes for the head dimension. For this, we first compute the weighting coefficients per head and average using them, followed by computing TokenRank for the aggregated result. We compare uniform weighting (simple average), random weighting, and $\lambda_{2}$ weighting. Results are presented in Table 6: While random weighting is significantly worse, $\lambda_{2}$ weighting performs better than simple averaging, particularly for transformers with a well-structured latent space, such as DINOv2 with registers.
Method uniform random $\lambda_{2}$
CLIP $\mathbf{49.07}$ $44.07$ $\mathbf{49.07}$
DINOv1 $53.50$ $47.21$ $\mathbf{53.55}$
DINOv2 $71.62$ $50.29$ $\mathbf{72.36}$
A.8 More SAG Results
In Figure 12 we show more examplary results comparing vanilla SD1.5 (Rombach et al., 2022), SAG (Hong et al., 2023) and SAG+TokenRank.
@@ A.8 More SAG Results
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
A.9 Per-Head TokenRank Visualizations
We show more visualizations for TokenRank, column sum, Center Patch, and CLS token for various heads and layers in Figure 15 and Figure 16.
@@ A.9 Per-Head TokenRank Visualizations
A.10 Visualization Across Generation Timesteps
In Figure 17, we show generated image results of using FLUX-Dev with 50 timesteps. We computed the TokenRank (incoming attention) and average over every 5 consecutive timesteps shown in every column (first entries correspond to noisier timesteps, while last entries are more noise-free). The TokenRank visualizations remain stable through the timestep dimension. It can be observed how attention maps are getting sharper during the denoising process, illustrating that TokenRank can serve to visualize and analyze generative models.

### v2

@@ E.2 $\lambda_{2}$ Weighting for Masking Most Influential Tokens Experiment
Appendix F Illustration of Bouncing as Consolidation Mechanism
Figure˜10 visualizes the first bounces and the steady state of two example images, heads and layers.
Figure˜11 also illustrates that iterating over the Markov chain results in an iterative refinement of the map, consolidating meta-stable states. The steady state eventually visualizes the global incoming attention. This effect cannot be observed for an input image with random pixels (Figure˜11, bottom). Instead, the Markov chain converges very quickly to a dispersed visualization after one step for a noisy input, as the attention matrix is not structured. We compute the visualizations by passing the images through DINOv1 and extracting the attention map in the 9th layer and 3rd head. Then, we compute the first bounce ($n=1$) with a uniform initial vector, the 2nd bounce, and the steady state. The same effect can be seen in Figure˜13, where the second bounce (Ours) is significantly cleaner and better adheres to the object boundaries compared to the first bounce (Column Select).
@@ Appendix F Illustration of Bouncing as Consolidation Mechanism
Appendix G More SAG Results
In Figure˜12 we show more results comparing vanilla SD1.5 (Rombach et al., 2022), SAG (Hong et al., 2023) and SAG+TokenRank.
@@ Appendix G More SAG Results
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Appendix H Per-Head TokenRank Visualizations
We show more visualizations for TokenRank, column sum, Center Patch, and CLS token for various heads and layers in Figure˜15 and Figure˜16.
@@ Appendix H Per-Head TokenRank Visualizations
Appendix I Visualization Across Generation Timesteps
In Figure˜17, we show generated image results of using FLUX-Dev with 50 timesteps. We computed the TokenRank (incoming attention) and average over every 5 consecutive timesteps shown in every column (first entries correspond to noisier timesteps, while last entries are more noise-free). The TokenRank visualizations remain stable through the timestep dimension. It can be observed how attention maps are getting sharper during the denoising process, illustrating that TokenRank can serve to visualize and analyze generative models.